OpenStudy (anonymous):

cos 5 theta + cos 3 theta =0

OpenStudy (linda3):

Welcome to OpenStudy @bbj2512b! :)

OpenStudy (solomonzelman):

\[\huge\color{blue}{ Cos(5θ) +Cos(3θ)=0 }\] Way 1: \[\color{blue}{ Cos(5θ) +Cos(3θ)=0 }\]\[\color{blue}{ Cos(5θ)=-Cos(3θ) }\]\[\color{blue}{ Cos(5θ) =Cos(3θ) }\]\[\color{blue}{ 5θ=3θ }\]\[So,~~~~θ~~~~is~~~~zero.\] Also (#2) : \[\color{blue}{ Cos(5θ)+Cos(3θ)=0 }\]\[\color{blue}{ Cos(4θ+θ)+Cos(4θ-0)=0 }\] See?

OpenStudy (solomonzelman):

The last equation in blue should be, \[Cos(4θ+θ)+Cos(4θ-θ)=0\]

OpenStudy (solomonzelman):

http://www.ies-math.com/math/java/trig/kahote/kahote.html

OpenStudy (solomonzelman):

I got disconnected.

OpenStudy (anonymous):

solve the equation by first transforming it into a product equal to 0 and then setting each factor equal to 0. use the following domain [0,360] in degrees or [0,2 pi] in radian cos 5 theta + cos 3 theta =0

OpenStudy (solomonzelman):

Oh, Idk how to do that, sorry :(

OpenStudy (solomonzelman):

Try tagging someone else like this, @bbj2512b

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

\[ -\cos(3\theta) \neq \cos(3\theta) \]And also, \(\theta = 0\) isn't even a solution since \(\cos(0) = 1\).